OpenStudy (anonymous):

?? With the restriction x ≠ 2, which of the following is the simplified form of 3 x plus 4 over x minus 2 minus the fraction x minus 8 over x minus 2 ? 1 4 times the quantity x plus 3 end quantity over x minus 2 x – 2 2 times the quantity x plus 6 end quantity over x minus 2

OpenStudy (anonymous):

\[\begin{align*}\frac{3x+4}{x-2}-\frac{x-8}{x-2}&=\frac{3x+4-(x-8)}{x-2}\\ &=\frac{2x+12}{x-2} \end{align*}\] How else can you write this?

OpenStudy (anonymous):

I'm not sure?

OpenStudy (anonymous):

@dreamer05 he's saying that you could factor the numerator by 2, making it\[\frac{ 2(x+6) }{ x-2 }\] making the answer 2 times the quantity x plus 6 end quantity over x minus 2